# Brain Mapping with Dense Features: Grounding Cortical Semantic Selectivity in Natural Images With Vision Transformers

### v1

@@ Abstract
Advances in large-scale artificial neural networks have facilitated novel insights into the functional topology of the brain. Here, we leverage this approach to study how semantic categories are organized in the human visual cortex. To overcome the challenge presented by the co-occurrence of multiple categories in natural images, we introduce BrainSAIL (Semantic Attribution and Image Localization), a method for isolating specific neurally-activating visual concepts in images. BrainSAIL exploits semantically consistent, dense spatial features from pre-trained vision models, building upon their demonstrated ability to robustly predict neural activity. This method derives clean, spatially dense embeddings without requiring any additional training, and employs a novel denoising process that leverages the semantic consistency of images under random augmentations. By unifying the space of whole-image embeddings and dense visual features and then applying voxel-wise encoding models to these features, we enable the identification of specific subregions of each image which drive selectivity patterns in different areas of the higher visual cortex. This provides a powerful tool for dissecting the neural mechanisms that underlie semantic visual processing for natural images. We validate BrainSAIL on cortical regions with known category selectivity, demonstrating its ability to accurately localize and disentangle selectivity to diverse visual concepts. Next, we demonstrate BrainSAIL’s ability to characterize high-level visual selectivity to scene properties and low-level visual features such as depth, luminance, and saturation, providing insights into the encoding of complex visual information. Finally, we use BrainSAIL to directly compare the feature selectivity of different brain encoding models across different regions of interest in visual cortex. Our innovative method paves the way for significant advances in mapping and decomposing high-level visual representations in the human brain.
@@ 1 Carnegie Mellon University 2 The University of Hong Kong
{afluo,jacobyeung,rzawar,srdewan,mmhender,lwehbe,michaeltarr}@cmu.edu footnotetext: * Co-corresponding authors.
@@ 2 Related Work
A growing body of work leveraging computational modeling and machine learning has explored semantic representation in the higher visual cortex. Approaches include generative image models (Ratan Murty et al., 2021; Gu et al., 2022; Pierzchlewicz et al., 2023; Luo et al., 2023; 2024) and the decoding of visual stimuli (Takagi & Nishimoto, 2022; Chen et al., 2022; Doerig et al., 2022; Ferrante et al., 2023; Liu et al., 2023; Scotti et al., 2024). These diverse studies are united by their consideration of the stimulus image as a whole, primarily focusing on the global information contained within the image rather than the individual scene components. In contrast, the method we introduce explicitly decomposes an image into its semantic components, enabling the identification of individual, semantically meaningful activating concepts within complex natural images.
@@
Using hand-crafted image stimuli, functional mapping studies have identified regions in the human brain that respond preferentially to stimuli representing distinct semantic concepts such as faces, places, bodies, words, objects, and food (Desimone et al., 1984; Sergent et al., 1992; Allison et al., 1994; McCarthy et al., 1997; Kanwisher et al., 1997; Gauthier & Tarr, 1997; Aguirre et al., 1996; Epstein & Kanwisher, 1998; Aguirre et al., 1998; O’Craven & Kanwisher, 2000; Nakamura et al., 2000; Aminoff et al., 2007; Downing et al., 2001; Cohen et al., 2000; Grill-Spector, 2003; Khosla et al., 2022; Pennock et al., 2023; Jain et al., 2023). One limitation of this simplified approach is that it may not fully capture the contextual complexity of natural vision (Gallant et al., 1998; Mahon, 2022). Addressing this concern, recent work on image-computable encoders has enabled computational tests of visual selectivity using naturalistic images  (Naselaris et al., 2011; Huth et al., 2012; Yamins et al., 2014; Eickenberg et al., 2017; Wen et al., 2018; Kubilius et al., 2019; Popham et al., 2021; Conwell et al., 2022; Wang et al., 2023a; Luo et al., 2023; Prince et al., 2023; Luo et al., 2024; Yang et al., 2024b; a; Efird et al., 2024). Building on this work, our method leverages state-of-the-art brain encoding backbones based on vision transformers (Dosovitskiy et al., 2020; Wang et al., 2023a) to further explore finer-grained semantic representation in visual cortex.
@@
There has been strong interest in leveraging generative models for decoding (reconstructing) visual stimuli conditioned on brain activations either directly or via intermediate language-based captions (Kamitani & Tong, 2005; Han et al., 2019; Seeliger et al., 2018; Shen et al., 2019; Ren et al., 2021; Takagi & Nishimoto, 2022; Chen et al., 2023; Lu et al., 2023; Ozcelik & VanRullen, 2023; Doerig et al., 2022; Ferrante et al., 2023; Liu et al., 2023; Mai & Zhang, 2023; Scotti et al., 2024). A related approach generates novel stimuli that are posited to best to activate a target brain region (as opposed to reconstructing the original stimulus) (Walker et al., 2019; Bashivan et al., 2019) with recent attempts utilizing GANs or Diffusion models to constrain the synthesized output (Ponce et al., 2019; Ratan Murty et al., 2021; Gu et al., 2022; Luo et al., 2023; 2024). While these models have shown positive results, they all rely on images as a whole, whereas BrainSAIL seeks to disentangle complex images into their semantically meaningful components and localize those parts of the image that elicit activation for different brain voxels or regions.
@@ 3 Methods
Figure 1: The BrainSAIL framework leverages dense visual features. (a) An fMRI encoder learns a map from images to voxel-wise activations in the brain. Encoders leveraging frozen foundation models based on vision transformers (ViTs) with voxel-wise adapters are currently the highest accuracy models for brain prediction (Conwell et al., 2022; Wang et al., 2023a). (b) Given an image and a ViT backbone for the fMRI encoder, we modify the backbone to output dense features. The dense backbone is wrapped inside of a Learning-Free Distillation Module. This module takes an image $\mathcal{I}$ and $2D$ image coordinates $\mathcal{C}$, and generates transformed images and coordinates $(\mathcal{I}_{i},\mathcal{C}_{i})$ for a given transform $\theta_{i}$. The dense features and transformed coordinates are provided to a denoising module to generate clean dense features. A voxel-wise adapter then generates dense relevance maps which highlight the image regions activating the voxel. (c) Using CLIP ViT-B/16 with the latest NACLIP adapter, we show relevance maps using the CLIP text encoder. The original, raw features are highly noisy and contain artifacts, while the distilled features are localized to the relevant semantic components with high accuracy. Note that we achieve state-of-the-art open vocabulary CLIP-based segmentation results using our method.
@@ 3.2 Deriving Dense Features from ViT backbones
MaskCLIP proposes to directly remove the convex re-weighting and output the value feature for each patch token directly. SCLIP and NACLIP reintroduce the weighting to reduce output artifacts, but modify it with correlative self-attention (CSA); or by using CSA with a spatial attentive bias $\omega$. Here, we utilize NACLIP as the dense adaptor for CLIP. The other two backbones in Section 4.4 use an updated ViT architecture with “register tokens” (Darcet et al., 2023). As these have not been explored in the context of CSA, we utilize MaskCLIP as the dense adaptor.
@@ 4.1 Setup
Figure 4: Grounding results using BrainSAIL. We visualize the top test set images as predicted by the CLIP fMRI encoder for each category selective region. For each image, we also visualize the image-wise UMAP for the distilled dense features. Note the UMAP basis here is computed imagewise, and not shared with Figure 3. For each image, we further visualize the feature relevancy map for the category selective voxels illustrating that this method extracts the semantically relevant regions in complex compositional images.
@@ 4.2 Image Factorization using the Brain
To explore how different areas in higher visual cortex align to different image parts we apply UMAP (McInnes et al., 2018) with an angular metric to linear brain weights and apply the same UMAP basis to dense features as produced by BrainSAIL. Note that during dimensionality reduction we do not utilize any cortex category masks from NSD – the region of interest outlines on the cortex in Figure 3a are for visualization purposes only and are derived from independent NSD functional localizers. As shown in Figure 3, we find that the factorization of the brain is well aligned to pre-identified functional regions, and broadly segments the cortex into axes along “people”, “scenes” and “food”. In particular, place regions, including the retrosplenial cortex (RSC), occipital place area (OPA), and parahippocampal place area (PPA), show selectivity for scene components (magenta). People regions, including the extrastriate body area (EBA), fusiform face area (FFA), occipital face area (OFA), show selectivity for face and body parts in the image (Green-Blue). Finally, we find that the recently identified food region that roughly surrounds FFA (YellowKhosla et al. (2022); Pennock et al. (2023); Jain et al. (2023) strongly corresponds to food in images. These results establish that BrainSAIL can be used to characterize higher-level selectivity to individual semantic categories in complex natural images without prior knowledge of their semantic selectivity.
We further quantify the feature relevance maps for broad category selective regions in Figure 4 and Table 2. We use the brain encoder to predict the top-5 images for the place/word/face/body regions, and the top-10 images for the food region. We find that our method can effectively localize the objects relevant to each category- selective brain region. Note that the word region is known to have cross-selectivity to faces (Mei et al., 2010) and food (Khosla & Wehbe, 2022).
@@ 4.3 Cortex Selectivity to Image Features
Going beyond semantic categories, we seek to explore the low- and mid-level image feature correlates that correspond to different brain regions. Prior work explored this by training a convolutional encoder on each NSD subject, which is limited to $\sim 10,000$ images each (Sarch et al., 2023). One concern is that using a small dataset with a convolutional backbone can lead to overfitting to the dataset’s specific features and exacerbate the inherent biases of convolutional networks. To address this limitation, our method leverages vision transformers trained on massive datasets of hundreds of millions of images, thereby avoiding the hard-coded inductive biases present in CNNs (Raghu et al., 2021). We visualize BrainSAIL feature dissection results in Figure 5. Our method can successfully identify the known scene selective regions (RSC/OPA/PPA) as preferring high depth, and is successful even in OPA where Sarch et al. (2023) fails. We believe this is likely because OPA processes higher-level associative content and affordances (Bonner & Epstein, 2017; Aminoff & Tarr, 2021). Similarly, we identify the region surrounding FFA as being selective to high color saturation, which correspond to the food regions identified by Jain et al. (2023) and others. In OPA, we identify a split in color luminance preference, which is similar to the indoor/outdoor preferring regions identified by Lescroart & Gallant (2019)Peer et al. (2019), and  Luo et al. (2023). These results demonstrate that our method can identify fine-grained selectivity with more broadly characterized brain regions.
@@ 4.4 Are Brain Encoders Equivalent?
To investigate the representational differences between the models, we perform BrainSAIL analysis for scene, face, and food-selective brain regions and qualitatively visualize the results in Figure 8. While all three models exhibit broad similarities in their grounding maps, DINO, trained without language supervision, demonstrates a stronger sensitivity to low-level visual features compared to CLIP and SigLIP (Wang et al., 2023a). This is evident in the food region (Figure 8), where DINO’s grounding map for a pizza image excludes the toppings and misses differently colored vegetables on a metal plate, suggesting a focus on color and texture rather than the concept of “food” itself. Similarly, in the face region, DINO’s grounding map exhibits less reliance on semantically relevant features such as eyes, nose, and mouth. We hypothesize that this greater sensitivity to visual features in DINO stems from its lack of language guidance during training, preventing it from learning the higher-level semantic correlations that link visually disparate parts and objects within a category. As high-performing “proxy models” of visual brain representation (Leeds et al., 2013), these and other underlying model characteristics – architecture, training objective, training dataset, etc. – are important considerations for developing more robust encoding models that can bridge the gap between artificial and biological vision systems.
@@ 5 Discussion
BrainSAIL achieves strong localization performance and benefits from a pre-trained vision transformer, reducing reliance on the fMRI dataset for backbone training. However, it is still necessary to train the fMRI encoder on these data, and thus potential dataset biases in the human neural data and how it was collected can influence the learned representations and conclusions. Future work should explore training on larger and more diverse neural datasets to mitigate this limitation and enhance the generalizability of our findings.
@@ Appendix A Appendix
  9. 9.
    Implementation details (A.9)
A.1 Additional encoder test set $R^{2}$ for CLIP, DINO, and SigLIP
A.2 Visualization of ground truth functional localizer statistics
A.3 Visualization of encoder weight UMAP for all subjects using CLIP backbone
A.4 Visualization of encoder weight UMAP for all subjects using DINO backbone
A.5 Visualization of encoder weight UMAP for all subjects using SigLIP backbone
A.6 Additional Visualization of Grounding using CLIP backbone
A.7 Additional Visualization of Grounding using DINO backbone
A.8 Additional Visualization of Grounding using SigLIP backbone
A.9 Implementation Details
@@
face_class = [”A face facing the camera”, ”A photo of a face”, ”A photo of a human face”, ”A photo of faces”, ”A photo of a person’s face”, ”A person looking at the camera”, ”People looking at the camera”,”A portrait of a person”, ”A portrait photo”] body_class = [”A photo of a torso”, ”A photo of torsos”, ”A photo of limbs”, ”A photo of bodies”, ”A photo of a person”, ”A photo of people”, ”A photo of a body”, ”A person’s arms”, ”A person’s legs”, ”A photo of hands”] scene_class = [”A photo of a bedroom”, ”A photo of an office”,”A photo of a hallway”, ”A photo of a doorway”, ”A photo of interior design”, ”A photo of a building”, ”A photo of a house”, ”A photo of nature”, ”A photo of landscape”, ”A landscape photo”, ”A photo of trees”, ”A photo of grass”] food_class = [”A photo of food”, ”A photo of cuisine”, ”A photo of fruit”, ”A photo of foodstuffs”, ”A photo of a meal”, ”A photo of bread”, ”A photo of rice”, ”A photo of a snack”, ”A photo of pastries”, ”A photo of vegetables”, ”A photo of pizza”, ”A photo of soup”, ”A photo of meat”, ”A photo of candy”] text_class = [”A photo of words”, ”A photo of glyphs”, ”A photo of a glyph”, ”A photo of text”, ”A photo of numbers”, ”A photo of a letter”, ”A photo of letters”, ”A photo of writing”, ”A photo of text on an object”]

### v2

@@ Abstract
We introduce BrainSAIL (Semantic Attribution and Image Localization), a method for linking neural selectivity with spatially distributed semantic visual concepts in natural scenes. BrainSAIL leverages recent advances in large-scale artificial neural networks, using them to provide insights into the functional topology of the brain. To overcome the challenge presented by the co-occurrence of multiple categories in natural images, BrainSAIL exploits semantically consistent, dense spatial features from pre-trained vision models, building upon their demonstrated ability to robustly predict neural activity. This method derives clean, spatially dense embeddings without requiring any additional training, and employs a novel denoising process that leverages the semantic consistency of images under random augmentations. By unifying the space of whole-image embeddings and dense visual features and then applying voxel-wise encoding models to these features, we enable the identification of specific subregions of each image which drive selectivity patterns in different areas of the higher visual cortex. This provides a powerful tool for dissecting the neural mechanisms that underlie semantic visual processing for natural images. We validate BrainSAIL on cortical regions with known category selectivity, demonstrating its ability to accurately localize and disentangle selectivity to diverse visual concepts. Next, we demonstrate BrainSAIL’s ability to characterize high-level visual selectivity to scene properties and low-level visual features such as depth, luminance, and saturation, providing insights into the encoding of complex visual information. Finally, we use BrainSAIL to directly compare the feature selectivity of different brain encoding models across different regions of interest in visual cortex. Our innovative method paves the way for significant advances in mapping and decomposing high-level visual representations in the human brain. Code available here.
@@ 1 Carnegie Mellon University 2 The University of Hong Kong
aluo@hku.hk, {jacobyeung,rzawar,srdewan,mmhender,
lwehbe,michaeltarr}@cmu.edu
footnotetext: * Co-corresponding authors.
@@ 2 Related Work
A growing body of work leveraging computational modeling and machine learning has explored semantic representation in the higher visual cortex. Approaches include generative image models (Ratan Murty et al., 2021; Gu et al., 2022; Pierzchlewicz et al., 2023; Luo et al., 2023; 2024) and the decoding of visual stimuli (Takagi & Nishimoto, 2022; Chen et al., 2022; Doerig et al., 2022; Ferrante et al., 2023; Liu et al., 2023; Scotti et al., 2024; Yeung et al., 2024). These diverse studies are united by their consideration of the stimulus image as a whole, primarily focusing on the global information contained within the image rather than the individual scene components. In contrast, the method we introduce decomposes an image into its semantic components, enabling the identification of individual, semantically meaningful activating concepts within complex natural images.
@@
Using hand-crafted image stimuli, functional mapping studies have identified regions in the human brain that respond preferentially to stimuli representing distinct semantic concepts such as faces, places, bodies, words, objects, and food (Desimone et al., 1984; Sergent et al., 1992; Allison et al., 1994; McCarthy et al., 1997; Kanwisher et al., 1997; Gauthier & Tarr, 1997; Aguirre et al., 1996; Epstein & Kanwisher, 1998; Aguirre et al., 1998; O’Craven & Kanwisher, 2000; Nakamura et al., 2000; Aminoff et al., 2007; Downing et al., 2001; Cohen et al., 2000; Grill-Spector, 2003; Khosla et al., 2022; Pennock et al., 2023; Jain et al., 2023). One limitation of this simplified approach is that it may not fully capture the contextual complexity of natural vision (Gallant et al., 1998; Mahon, 2022). Addressing this concern, recent work on image-computable encoders has enabled computational tests of visual selectivity using naturalistic images  (Naselaris et al., 2011; Huth et al., 2012; Yamins et al., 2014; Eickenberg et al., 2017; Wen et al., 2018; Kubilius et al., 2019; Popham et al., 2021; Conwell et al., 2022; Wang et al., 2023a; Luo et al., 2023; Prince et al., 2023; Adeli et al., 2023; Luo et al., 2024; Yang et al., 2024b; a; Efird et al., 2024). Building on this work, our method leverages state-of-the-art brain encoding backbones based on vision transformers (Dosovitskiy et al., 2020; Wang et al., 2023a) to further explore finer-grained semantic representation in visual cortex.
@@
There has been strong interest in leveraging generative models for decoding (reconstructing) visual stimuli conditioned on brain activations (Kamitani & Tong, 2005; Han et al., 2019; Seeliger et al., 2018; Shen et al., 2019; Ren et al., 2021; Takagi & Nishimoto, 2022; Chen et al., 2023; Lu et al., 2023; Ozcelik & VanRullen, 2023; Doerig et al., 2022; Ferrante et al., 2023; Liu et al., 2023; Mai & Zhang, 2023; Scotti et al., 2024). A related approach generates novel stimuli that are posited to best to activate a target brain region (as opposed to reconstructing the original stimulus) (Walker et al., 2019; Bashivan et al., 2019) with recent attempts utilizing GANs or Diffusion models to constrain the synthesized output (Ponce et al., 2019; Ratan Murty et al., 2021; Gu et al., 2022; Luo et al., 2023; 2024). While these models have shown positive results, they all rely on images as a whole, whereas BrainSAIL seeks to disentangle complex images into their semantically meaningful components and localize those parts of the image that elicit activation for different brain voxels or regions.
@@ 3 Methods
Figure 1: The BrainSAIL framework leverages dense visual features. (a) An fMRI encoder learns a map from images to voxel-wise activations in the brain. Encoders leveraging frozen foundation models based on vision transformers (ViTs) with voxel-wise adapters are currently the highest accuracy models for brain prediction (Conwell et al., 2022; Wang et al., 2023a). (b) Given an image and a ViT backbone for the fMRI encoder, we modify the backbone to output dense features. The dense backbone is wrapped inside of a Learning-Free Distillation Module. This module takes an image $\mathcal{I}$ and $2D$ image coordinates $\mathcal{C}$, and generates transformed images and coordinates $(\mathcal{I}_{i},\mathcal{C}_{i})$ for a given transform $\theta_{i}$. The dense features and transformed coordinates are provided to a denoising module to generate clean dense features. The frozen voxel-wise adapter from (a) is applied to each patch to generate dense relevance maps which highlight the image regions activating the voxel. (c) Using CLIP ViT-B/16 with the latest NACLIP adapter, we show relevance maps using the CLIP text encoder. The NACLIP raw features are highly noisy and contain artifacts, while the distilled features are localized to the relevant semantic components with high accuracy. Note that we achieve state-of-the-art open vocabulary CLIP-based segmentation results using our method.
@@ 3.2 Deriving Dense Features from ViT backbones
MaskCLIP directly removes the convex re-weighting and outputs the value feature for each patch token directly. SCLIP and NACLIP reintroduce the weighting to reduce output artifacts, but modify it with correlative self-attention (CSA); or by using CSA with a spatial attentive bias $\omega$. Here, we utilize NACLIP as the dense adaptor for CLIP. The other two backbones in Section 4.4 use an updated ViT architecture with “register tokens” (Darcet et al., 2023). As these have not been explored in the context of CSA, we utilize MaskCLIP as the dense adaptor. We treat dense features as unit-norm.
@@ 4.1 Setup
Figure 4: Grounding results using BrainSAIL. We visualize the top test set images as predicted by the CLIP fMRI encoder for each category selective region. For each image, we also visualize the image-wise UMAP for the distilled dense features. Note the UMAP basis here is computed imagewise, and not shared with Figure 3. The image-wise UMAP shows the semantic components present in the dense features. For each image, we further visualize the feature relevancy map for the category selective voxels illustrating that this method extracts the semantically relevant regions in complex compositional images.
@@ 4.2 Image Factorization using the Brain
To explore how different category-responsive regions in higher visual cortex align to different image parts we apply UMAP (McInnes et al., 2018) with an angular metric to linear brain weights and apply the same UMAP basis to dense features as produced by BrainSAIL. Note that during dimensionality reduction we do not utilize any cortex category masks from NSD – the region of interest outlines on the cortex in Figure 3a are for visualization purposes only and are derived from independent NSD functional localizers. As shown in Figure 3, we find that the factorization of the brain is well aligned to pre-identified functional regions, and broadly segments the cortex into axes along “people”, “scenes” and “food”. In particular, place regions, including the retrosplenial cortex (RSC), occipital place area (OPA), and parahippocampal place area (PPA), show selectivity for scene components (magenta). People regions, including the extrastriate body area (EBA), fusiform face area (FFA), occipital face area (OFA), show selectivity for face and body parts in the image (Green-Blue). Finally, we find that the recently identified food-responsive region that roughly surrounds FFA (YellowKhosla et al. (2022); Pennock et al. (2023); Jain et al. (2023) strongly corresponds to food in images. These results establish that BrainSAIL can be used to characterize higher-level selectivity to individual semantic categories in complex natural images without prior knowledge of their semantic selectivity.
We further quantify the feature relevance maps for broad category selective regions in Figure 4 and Table 2. We use the brain encoder to predict the top-5 images for the place/word/face/body regions, and the top-10 images for the food-responsive region. We find that our method can effectively localize the objects relevant to each category- selective brain region. Note that the word region is known to have cross-selectivity to faces (Mei et al., 2010) and food (Khosla & Wehbe, 2022).
@@ 4.3 Cortex Selectivity to Image Features
Going beyond semantic categories, we seek to explore the low- and mid-level image feature correlates that correspond to different brain regions. Prior work explored this by training a convolutional encoder on each NSD subject, which is limited to $\sim 10,000$ images each (Sarch et al., 2023). One concern is that using a small dataset with a convolutional backbone can lead to overfitting to the dataset’s specific features and exacerbate the inherent biases of convolutional networks. To address this limitation, our method leverages vision transformers trained on massive datasets of hundreds of millions of images, thereby avoiding the hard-coded inductive biases present in CNNs (Raghu et al., 2021). We visualize BrainSAIL feature dissection results in Figure 5. Our method can successfully identify the known scene selective regions (RSC/OPA/PPA) as preferring high depth, and is successful even in OPA where Sarch et al. (2023) fails. We believe this is likely because OPA processes higher-level associative content and affordances (Bonner & Epstein, 2017; Aminoff & Tarr, 2021). Similarly, we identify the region surrounding FFA as being selective to high color saturation, which correspond to the food-responsive regions identified by Jain et al. (2023) and others. In OPA, we identify a split in color luminance preference, which is similar to the indoor/outdoor preferring regions identified by Lescroart & Gallant (2019)Peer et al. (2019), and  Luo et al. (2023). These results demonstrate that our method can identify fine-grained selectivity with more broadly characterized brain regions.
@@ 4.4 Are Brain Encoders Equivalent?
To investigate the representational differences between the models, we perform BrainSAIL analysis for scene-, face-, and food-responsive brain regions and qualitatively visualize the results in Figure 8. While all three models exhibit broad similarities in their grounding maps, DINO, trained without language supervision, demonstrates a stronger sensitivity to low-level visual features compared to CLIP and SigLIP (Wang et al., 2023a). This is evident in the food-responsive region (Figure 8), where DINO’s grounding map for a pizza image excludes the toppings and misses differently colored vegetables on a metal plate, suggesting a focus on color and texture rather than the concept of “food” itself. Similarly, in the face region, DINO’s grounding map exhibits less reliance on semantically relevant features such as eyes, nose, and mouth. We hypothesize that this greater sensitivity to visual features in DINO stems from its lack of language guidance during training, preventing it from learning the higher-level semantic correlations that link visually disparate parts and objects within a category. As high-performing “proxy models” of visual brain representation (Leeds et al., 2013), these and other underlying model characteristics – architecture, training objective, training dataset, etc. – are important considerations for developing more robust encoding models that can bridge the gap between artificial and biological vision systems.
@@ 5 Discussion
BrainSAIL achieves strong localization performance and benefits from a pre-trained vision transformer, reducing reliance on the fMRI dataset for backbone training. However, it is still necessary to train the fMRI encoder on these data, and thus potential dataset biases in the human neural data and how it was collected can influence the learned representations and conclusions (Shirakawa et al., 2024). Future work should explore training on larger and more diverse neural datasets to mitigate this limitation and enhance the generalizability of our findings.
@@ Appendix A Appendix
    Implementation and fMRI processing details (A.1)
  2. 2.
    Comparing CLIP granularity and brain saliency (A.2)
  3. 3.
@@
A.1 Implementation and fMRI Processing Details
@@
face_class = [”A face facing the camera”, ”A photo of a face”, ”A photo of a human face”, ”A photo of faces”, ”A photo of a person’s face”, ”A person looking at the camera”, ”People looking at the camera”,”A portrait of a person”, ”A portrait photo”]
body_class = [”A photo of a torso”, ”A photo of torsos”, ”A photo of limbs”, ”A photo of bodies”, ”A photo of a person”, ”A photo of people”, ”A photo of a body”, ”A person’s arms”, ”A person’s legs”, ”A photo of hands”]
scene_class = [”A photo of a bedroom”, ”A photo of an office”,”A photo of a hallway”, ”A photo of a doorway”, ”A photo of interior design”, ”A photo of a building”, ”A photo of a house”, ”A photo of nature”, ”A photo of landscape”, ”A landscape photo”, ”A photo of trees”, ”A photo of grass”]
food_class = [”A photo of food”, ”A photo of cuisine”, ”A photo of fruit”, ”A photo of foodstuffs”, ”A photo of a meal”, ”A photo of bread”, ”A photo of rice”, ”A photo of a snack”, ”A photo of pastries”, ”A photo of vegetables”, ”A photo of pizza”, ”A photo of soup”, ”A photo of meat”, ”A photo of candy”]
text_class = [”A photo of words”, ”A photo of glyphs”, ”A photo of a glyph”, ”A photo of text”, ”A photo of numbers”, ”A photo of a letter”, ”A photo of letters”, ”A photo of writing”, ”A photo of text on an object”]
fMRI data processing.
All of our analysis is done in subject native space voxels in $1.8mm^{3}$ resolution (func1pt8mm), and we use the beta values from betas_fithrf_GLMdenoise_RR. Cortical voxels are selected based on the nsdgeneral mask.
In order to select visually responsive voxels for Figure 3, we utilize the HCP parcellation which yields 180 regions. For each region in each subject, we compute the average noise ceiling using NSD provided data. We rank the regions within each subject, and average the ranks across all subjects. The best $25\%$ (45 out of 180) regions by noise ceiling are selected. We further mask out voxels which are labeled as early visual (V1$\sim$ V4).
Intuition for learning-free distillation.
We are specifically motivated by literature in learning dense semantic embeddings. In particular,  Kobayashi et al. (2022) proposed learning a dense 3D semantic representation. However Kobayashi et al. (2022) uses LSeg (Li et al., 2022) – a supervised open-vocabulary segmentation network which demonstrates relatively poor generalization to categories outside of the training set (Kerr et al., 2023). As we do not wish to restrict the hypothesis space relative to cortical selectivity, we were motivated to explore methods which preserve the original representational capability of vision foundation models. By default, the dense token embeddings do not lie in the same semantic space as the summary token. However modifications like MaskCLIP (Zhou et al., 2022), SCLIP (Wang et al., 2023b), and NACLIP (Hajimiri et al., 2024) propose changes to the self-attention mechanism that ensure the dense embeddings are in the same space as the summary token. We utilize NACLIP for CLIP (or models without registers). As register tokens have not been explored in the context of self-attention modifications, we use MaskCLIP for models with registers. Without further distillation, these methods yield dense embeddings that have artifacts.
Recent work has proposed to learn consistent 3D representations using NeRF and feature fields (Mildenhall et al., 2021; Kobayashi et al., 2022). These methods demonstrate that a spatially consistent representation can be learned in the presence of noise (pixel noise or measurement noise) by utilizing a photometric 3D consistency framework with multiple views. As we do not have multiple views of a single image in 2D, we augment images with horizontal flips and vertical/horizontal offsets to generate synthetic views. These augmentations do not change the semantics of a scene. We observe that because the artifacts are not static relative to scene content, augmentations can effectively render artifacts into ”measurement noise”.
Originally NeRF and feature fields perform very expensive neural field learning with gradients and MSE loss. Faster learning and inference procedures are required to facilitate data-driven studies of the visual cortex using thousands of images. We show that averaging over views achieves the same mathematical result as MSE driven optimization.
A.2 Comparing CLIP granularity and brain saliency
Figure S.1: Comparing the granularity of CLIP embeddings with brain region saliency. On the left, we visualize the image and the UMAP of the distilled NACLIP embeddings. On the top right, we probe the dense features using the CLIP-text encoder. On the bottom right, we visualize the relevance map using the fMRI encoder weights for $t>2$ functional regions from S1. In both cases, the probing is done with a cosine similarity. We find that brain data usually has semantically coarser attributions. Likely due to NSD data co-occurrence statistics, we find that neural data does not consistently separate faces & bodies, while CLIP itself can have very fine-grained representations.
Figure S.2: Comparing the granularity of CLIP embeddings with brain region saliency. On the left, we visualize the image and the UMAP of the distilled NACLIP embeddings. On the top right, we probe the dense features using the CLIP-text encoder. On the bottom right, we visualize the relevance map using the fMRI encoder weights for $t>2$ functional regions from S1. In both cases, the probing is done with a cosine similarity. We find that brain data usually has semantically coarser attributions. Likely due to NSD data co-occurrence statistics, we find that neural data does not consistently separate faces & bodies, and attributes food selectivity to small objects or high saturation image regions (note sunglasses in top image, and pink dress of bottom image) – food itself is often both small and high saturation. Our method can reveal surprising behavior of fMRI encoders.
A.3 Additional encoder test set $R^{2}$ for CLIP, DINO, and SigLIP
A.4 Visualization of ground truth functional localizer statistics
A.5 Visualization of encoder weight UMAP for all subjects using CLIP backbone
A.6 Visualization of encoder weight UMAP for all subjects using DINO backbone
A.7 Visualization of encoder weight UMAP for all subjects using SigLIP backbone
A.8 Additional Visualization of Grounding using CLIP backbone
A.9 Additional Visualization of Grounding using DINO backbone
A.10 Additional Visualization of Grounding using SigLIP backbone